Quantities in Chemical Reactions

  1. Ricardo states that for the process of iron reacting with sulfur to produce iron (III) sulfide according to

16 Fe(s) + 3 S8(s) → 8 Fe2S3(s),

16.0 g of iron will react with 3.00 g of sulfur in the reaction to produce 8.00g of iron (III) sulfide. Is Ricardo's

reasoning correct? Explain your answer.


2. A magnesium compound is made up of magnesium and another element. It contains 72.2% magnesium by mass

and the rest is made up by the other element. If the molar mass of the compound is 100.95 g/mol, identify the

other element.


3. Sodium hydroxide is commonly called caustic soda, and is used in many household drain cleaners. It is prepared

by the reaction:

calcium hydroxide + sodium carbonate → sodium hydroxide + calcium carbonate

a) Write the balanced equation for the reaction.


b) What is the maximum mass of sodium hydroxide that could be produced from 195 g of calcium hydroxide?


1
Expert's answer
2021-03-31T07:57:28-0400

1 . Ricardo reasoning is wrong.

The limiting reagent is S8

so 3S8 =8Fe2S3

3g Sul[hur = (8 x 88 x 3)/(3 x 256.5) Fe2S3

= 2.74g Fe2S3


2 . % of element = 100-72.2 = 27.8

27.8% x 100.95 = 28.06

since the valency of magnesium is =2 , then the atomic mass will be 28/2 =14 which is nitrogen gas

therefore, the formula of the compound is Mg3N2


3 a . Ca(OH)2 + Na2CO3 = 2NaOH + CaCO3

b. Mass of NaOH = (2 x 40 x 195)/ 74 = 210.81 g NaOH